# Bounds on the number of squares in recurrence sequences

### v2

@@ 4.1. Prerequisites
We will suppose that there are two distinct squares, $y_{k}$ and $y_{\ell}$, in the sequence with $y_{\ell}>y_{k}>1$. So $y_{\ell}>y_{k}\geq 4\geq\max\left\{4\sqrt{\left|N_{\alpha}\right|/d},\left|N_{\alpha}\right|/d\right\}$ (since $b=1$ and $N_{\alpha}=a^{2}-bd^{4}<0$, so $\left|N_{\alpha}\right|<d$), as required in our lemmas above.
@@
As in Lemma 3.9, we put $\omega=\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)/\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)$ and let $\zeta_{4}$ be the $4$-th root of unity such that
$\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|=\min_{0\leq j\leq 3}\left|\omega^{1/4}-e^{2j\pi i/4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|,$
where $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}=\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$ with $r_{k}$, $r_{\ell}$, $s_{k}$ and $s_{\ell}$ as in Proposition 3.1, which are associated with $\left(x_{k},y_{k}\right)$ and $\left(x_{\ell},y_{\ell}\right)$, respectively. From (3.31) in the proof of Lemma 3.9(b), we have
$\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.127.$
@@
(4.2) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}=\left|\omega-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|>3.959\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|.$
@@
By our choice of $k$ and $\ell$, and by Lemma 3.9(b) when $-N_{\alpha}$ is not a square, $\zeta_{4}=\pm 1\in\mathbb{Q}\left(\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$. This is important for us here as $\zeta_{4}\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)/\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$ must be in an imaginary quadratic field in order to apply Lemma 2.1 to obtain a lower bound for the rightmost quantity in (4.2).
@@
Writing $\omega=e^{i\varphi}$, with $-\pi<\varphi\leq\pi$, from Lemmas 2.3(a) and 3.9(a), we can take
(4.8) $\ell_{0}=\mathcal{C}_{4,2}|\varphi|<0.458\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|.$
Also from Lemma 3.9(a), we have $|\varphi|<0.6$, so the condition $|\omega-1|<1$ in Lemma 2.2 is satisfied.
@@
We are now ready to deduce the required contradiction from the assumption that there are two sufficiently large squares when $-N_{\alpha}$ is a square. We will break the proof into five parts.
@@ 4.2. $r_{0}=1$ and $\zeta_{4}p/q\neq p_{1}/q_{1}$ for all $4$-th roots of unity, $\zeta_{4}$
From the equality in (4.2), along with Lemma 2.1(b) and (4.9), we have
(4.10) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>3.959\left|\omega^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|>\frac{3.959(1-c)}{k_{0}Q^{r_{0}}\sqrt{f_{k}f_{\ell}}\left(y_{k}y_{\ell}\right)^{1/4}}.$
@@
We will now combine (4.12) with the gap principle in Lemma 3.8(a). to show that this case cannot occur.
Since $-N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b), so combining the upper bound for $y_{\ell}^{3}$ in (4.12) with Lemma 3.8(a) and cancelling the common factor of $y_{k}^{5}$ on both sides, we find that
@@
By (4.1), we have $0.250.24/\left((1-c)d^{1/2}\right)<$ 0.25 if $c<0.9$. We will see in Subsection 4.4 that $c=0.75$ is a good choice. So $y_{k}<\left|N_{\alpha}\right|/4$, which contradicts parts (b) and (c) of Lemma 3.5. Hence this case cannot hold.
@@ 4.3. $r_{0}=1$ and $\zeta_{4}p/q=p_{1}/q_{1}$ for some $4$-th root of unity, $\zeta_{4}$
$\left|\omega^{1/4}-\zeta_{4}\frac{p}{q}\right|=\frac{1}{q_{r_{0}}}\left|q_{r_{0}}\omega^{1/4}-p_{r_{0}}\right|$
$=$ $\left|\frac{N_{d^{\prime},4,r_{0}}}{D_{4,r_{0}}Y_{1,4,r_{0}}(\omega)\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r_{0}}}\right|\left|\frac{D_{4,r_{0}}}{N_{d^{\prime},4,r_{0}}}R_{1,4,r_{0}}(\omega)\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r_{0}}\right|$
$=$ $\left|\frac{(\omega-1)^{2r_{0}+1}}{Y_{1,4,r_{0}}(\omega)}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}{}_{2}F_{1}\left(r_{0}+3/4,r_{0}+1;2r_{0}+2;1-\omega\right)\right|$
(4.13) $\geq$ $\left|\frac{(\omega-1)^{2r_{0}+1}}{Y_{1,4,r_{0}}(\omega)}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}\right|.$
@@
(4.14) $\left|Y_{1,4,r_{0}}(\omega)\right|<1.072\frac{r_{0}!\Gamma(3/4)}{\Gamma(r_{0}+3/4)}\left|1+\sqrt{\omega}\right|^{2r_{0}}.$
We can write $\omega-1$ as $\left(2N_{\alpha}+2x_{k}N_{\varepsilon}\sqrt{N_{\alpha}}\right)/\left(x_{k}^{2}-N_{\alpha}\right)$, so
$|\omega-1|=\sqrt{\frac{4\left|N_{\alpha}\right|}{x_{k}^{2}-N_{\alpha}}}=2\sqrt{\frac{\left|N_{\alpha}\right|}{d}}\frac{1}{y_{k}}.$
Since $\left|\sqrt{\omega}+1\right|<2$, it follows that
$\left|\sqrt{\omega}-1\right|=\frac{\left|\omega-1\right|}{\left|\sqrt{\omega}+1\right|}>\sqrt{\frac{\left|N_{\alpha}\right|}{d}}\frac{1}{y_{k}}.$
@@
$\left|\frac{(\omega-1)^{2r_{0}+1}}{\left(1+\sqrt{\omega}\right)^{2r_{0}}}\right|=\left|(\omega-1)\left(\sqrt{\omega}-1\right)^{2r_{0}}\right|>2\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1}.$
Applying this inequality together with the upper bound for $\left|Y_{1,4,r_{0}}(\omega)\right|$ in (4.14) and the inequalities (2.3) in Lemma 2.3(b) to (4.3), it follows that
$\left|\omega^{1/4}-\zeta_{4}\frac{p}{q}\right|$ $>\frac{0.2915}{4^{r_{0}}\cdot r_{0}^{1/2}}\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1}.$
@@
since $-N_{\alpha}$ is a square and $d<\left|N_{\alpha}\right|$. But this inequality is impossible. Hence this case cannot hold.
4.4. $r_{0}>1$, $\zeta_{4}p/q\neq p_{r_{0}}/q_{r_{0}}$ for all $4$-th roots of unity, $\zeta_{4}$
@@
Since $-N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b), so
@@ 4.5. $r_{0}>1$ and $\zeta_{4}p/q=p_{r_{0}}/q_{r_{0}}$ for some $4$-th root of unity, $\zeta_{4}$
Since $-N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b). Also, as $r_{0}-1\geq 1$, it follows that $0.7405\cdot 0.0002344^{r_{0}-1}>0.0001735^{r_{0}-1}$. So the above inequality implies
@@ 4.6. Small $d$
We check directly all pairs $(a,d)$ of positive integers with $2\leq d\leq 104$ not a square and $-N_{\alpha}=d-a^{2}$ a square. Note that there are only finitely many such pairs, since $N_{\alpha}=a^{2}-d<0$.
@@
Next we treat the case when $K<-1$, where, as in Lemma 3.5, $K$ is the largest negative integer such that $y_{K}>1$.
@@
If $K<-1$, then there are at most two distinct integer squares among the $y_{k}$’s.
@@
From Lemma 3.6, we know that $K<-1$ can only happen in the following two cases:
@@
For the remainder of this section, we may assume that $K=-1$.
@@
Similarly, if $-N_{\alpha}$ is an odd square with $N_{\alpha}\leq-9$ and $u\geq 3$, then we have $y_{k}>(81/10)\left|N_{\alpha}\right|^{2}\geq(729/10)\left|N_{\alpha}\right|>(729/10)\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$. But, we saw at the start of this section that for $N_{\alpha}$ odd, there can be no squares $y_{k}$ and $y_{\ell}$ satisfying $y_{\ell}>y_{k}>\max\left(1,19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}\right)$. So we need only consider $u\leq 2$ when $N_{\alpha}$ is odd.
@@
For $u=1$, we proceed similarly. Here we have $y_{k}>\left|N_{\alpha}\right|^{2}/10$ for $|k|\geq 2$, so we need to consider $4\leq\left|N_{\alpha}\right|\leq 36$ when $-N_{\alpha}$ is an even square and $9\leq\left|N_{\alpha}\right|\leq 169$ when $-N_{\alpha}$ is an odd square.
@@
$y_{1}>\frac{57.32d^{2}}{\left|N_{\alpha}\right|^{2}}y_{-1}\geq\frac{57.32d^{2}}{\left|N_{\alpha}\right|^{2}}\left(\frac{\left|N_{\alpha}\right|u^{2}}{4}\right)^{3}=\frac{57.32d^{2}\left|N_{\alpha}\right|u^{6}}{64}.$
@@
Suppose that $-N_{\alpha}$ is an odd square and $y_{\pm 1}$ is a square integer.
@@
(a) Next we show that if $t^{2}-du^{2}=4$, then $4\nmid t$. If $4\mid t$, then $du^{2}\equiv 12\pmod{16}$. Since $-N_{\alpha}$ is an odd square, we have $-N_{\alpha}\equiv 1,9\pmod{16}$. If $a$ is odd, then $a^{2}\equiv 1\pmod{8}$. So $d\equiv 2\pmod{8}$ and hence $u^{2}\equiv 6\pmod{8}$, which is not possible. If $a$ is even, then $a^{2}\equiv 0,4\pmod{16}$. Since $-N_{\alpha}=d-a^{2}\equiv 1,9\pmod{16}$, it follows that $d\equiv 1\pmod{8}$. So $u^{2}\equiv 12\pmod{16}$, which is also impossible. So $4\nmid t$.
@@
Since $N_{\alpha}$ is odd, we have $-N_{\alpha}\equiv 1,9\pmod{16}$. Also $a^{2}\equiv 0,1,4,9\pmod{16}$. Hence $d\equiv 1,2,5,9,10,13\pmod{16}$ and so $8|u^{2}$. Thus $4|u$.
@@
If $-N_{\alpha}$ is an odd square, $1\leq u\leq 8$ and $y_{\pm 1}\in\mathbb{Z}$ a square with $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $u=2$, $t^{2}-du^{2}=-4$ and $\gcd\left(a^{2},d\right)=1$.
@@
From Lemma 5.4(a), if $t^{2}-du^{2}=4$ and $-N_{\alpha}$ is an odd square, then $4|u$, so we are left with $u=4$ or $8$ here.
@@
Arguing modulo $9$, from $t^{2}-64d=4$, we see that $d\equiv 0,3,5,6\pmod{9}$. From this, $-N_{\alpha}=d-a^{2}$ being a square and the squares modulo $9$ being $0,1,4,7$, we must have $\left(a^{2}\bmod 9,d\bmod 9\right)=(0,0)$, $(1,5)$, $(4,5)$ or $(7,5)$.
@@
Therefore by Lemma 3.12, we have $\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\geq 36$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}<(76/36)\left|N_{\alpha}\right|$. But by Lemma 3.5(c), we know that $y_{\pm 1}\geq 16\left|N_{\alpha}\right|$ here. So the case where $u=8$, $-N_{\alpha}$ is an odd square and $t^{2}-du^{2}=4$ is excluded.
@@
We suppose otherwise. Since $N_{\alpha}$ is odd, $\gcd\left(a^{2},d\right)$ must be odd and since $d=\left(t^{2}+4\right)/4$, any odd prime factor, $p$, of the gcd must also be a factor of $t^{2}+4$. This means that $t^{2}\equiv-4\pmod{p}$. Hence $p\equiv 1\pmod{4}$. Since $p^{2}|a^{2}$ and $-N_{\alpha}=d-a^{2}$ is a perfect square, this also implies that $p^{2}|d$. Therefore by Lemma 3.12, we have $\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\geq 100$. Hence the theorem holds if $y_{\pm 1}>(76/100)\left|N_{\alpha}\right|$, by Proposition 4.1. But by Lemma 3.5(c), we know that $y_{\pm 1}\geq\left|N_{\alpha}\right|$ here. So the theorem holds if $u=2$, $-N_{\alpha}$ is an odd square and $\gcd\left(a^{2},d\right)>1$. ∎

### v3

@@ 4.1. Prerequisites
We will suppose that there are two distinct squares, $y_{k}$ and $y_{\ell}$, in the sequence with $y_{\ell}>y_{k}>1$. So $y_{\ell}>y_{k}\geq 4\geq\max\left(4\sqrt{\left|N_{\alpha}\right|/d},\left|N_{\alpha}\right|/d\right)$ (since $b=1$ and $N_{\alpha}=a^{2}-bd^{4}<0$, so $\left|N_{\alpha}\right|<d$), as required in our lemmas above.
@@
As in Lemma 3.9, we put $\omega_{k}=\left(x_{k}+N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)/\left(x_{k}-N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)$ and let $\zeta_{4}$ be the $4$-th root of unity such that
$\left|\omega_{k}^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|=\min_{0\leq j\leq 3}\left|\omega_{k}^{1/4}-e^{2j\pi i/4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|,$
where $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}=\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$ with $\left(r_{k},s_{k}\right)$ and $\left(r_{\ell},s_{\ell}\right)$ as in Proposition 3.1, which are associated with $\left(x_{k},y_{k}\right)$ and $\left(x_{\ell},y_{\ell}\right)$, respectively. From (3.31) in the proof of Lemma 3.9(b), we have
$\left|\omega_{k}^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|<0.127.$
@@
(4.2) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}=\left|\omega_{k}-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|>3.959\left|\omega_{k}^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|.$
@@
By our choice of $k$ and $\ell$, and by Lemma 3.9(b) when $N_{\alpha}$ is not a square, $\zeta_{4}=\pm 1\in\mathbb{Q}\left(\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$. This is important for us here as $\zeta_{4}\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)/\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$ must be in an imaginary quadratic field in order to apply Lemma 2.1 to obtain a lower bound for the rightmost quantity in (4.2).
@@
Writing $\omega_{k}=e^{i\varphi_{k}}$, with $\pi<\varphi_{k}\leq\pi$, from Lemmas 2.3(a) and 3.9(a), we can take
(4.8) $\ell_{0}=\mathcal{C}_{4,2}\left|\varphi_{k}\right|<0.458\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|.$
Also from Lemma 3.9(a), we have $\left|\varphi_{k}\right|<0.6$, so the condition $\left|\omega_{k}-1\right|<1$ in Lemma 2.2 is satisfied.
@@
We are now ready to deduce the required contradiction from the assumption that there are two sufficiently large squares when $N_{\alpha}$ is a square. We will break the proof into five parts.
@@ 4.2. $r_{0}=1$ and $\zeta_{4}p/q\neq p_{1}/q_{1}$ for all $4$-th roots of unity, $\zeta_{4}$
From (4.2), along with Lemma 2.1(b) and (4.9), we have
(4.10) $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}>3.959\left|\omega_{k}^{1/4}-\zeta_{4}\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|>\frac{3.959(1-c)}{k_{0}Q^{r_{0}}\sqrt{f_{k}f_{\ell}}\left(y_{k}y_{\ell}\right)^{1/4}}.$
@@
We will now combine (4.12) with the gap principle in Lemma 3.8(a) to show that this case cannot occur.
Since $N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b), so combining the upper bound for $y_{\ell}^{3}$ in (4.12) with Lemma 3.8(a) and cancelling the common factor of $y_{k}^{5}$ on both sides, we find that
@@
By (4.1), we have $0.24/\left((1-c)d^{1/2}\right)<0.25$ if $c<0.9$. We will see in Subsection 4.4 that $c=0.75$ is a good choice. So $y_{k}<\left|N_{\alpha}\right|/4$, which contradicts parts (b) and (c) of Lemma 3.5. Hence this case cannot hold.
@@ 4.3. $r_{0}=1$ and $\zeta_{4}p/q=p_{1}/q_{1}$ for some $4$-th root of unity, $\zeta_{4}$
$\left|\omega_{k}^{1/4}-\zeta_{4}\frac{p}{q}\right|=\frac{1}{q_{r_{0}}}\left|q_{r_{0}}\omega_{k}^{1/4}-p_{r_{0}}\right|$
$=$ $\left|\frac{N_{d^{\prime},4,r_{0}}}{D_{4,r_{0}}Y_{1,4,r_{0}}(\omega_{k})\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r_{0}}}\right|\left|\frac{D_{4,r_{0}}}{N_{d^{\prime},4,r_{0}}}R_{1,4,r_{0}}(\omega_{k})\left(\frac{u_{1}-u_{2}\sqrt{t^{\prime}}}{2g}\right)^{r_{0}}\right|$
$=$ $\left|\frac{(\omega_{k}-1)^{2r_{0}+1}}{Y_{1,4,r_{0}}(\omega_{k})}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}{}_{2}F_{1}\left(r_{0}+3/4,r_{0}+1;2r_{0}+2;1-\omega_{k}\right)\right|$
(4.13) $\geq$ $\left|\frac{(\omega_{k}-1)^{2r_{0}+1}}{Y_{1,4,r_{0}}(\omega_{k})}\frac{(1/4)\cdots(r_{0}+1/4)}{(r_{0}+1)\cdots(2r_{0}+1)}\right|.$
@@
(4.14) $\left|Y_{1,4,r_{0}}\left(\omega_{k}\right)\right|<1.072\frac{r_{0}!\Gamma(3/4)}{\Gamma(r_{0}+3/4)}\left|1+\sqrt{\omega_{k}}\right|^{2r_{0}}.$
We can write $\omega_{k}-1$ as $\left(2N_{\alpha}+2x_{k}N_{\varepsilon^{k}}\sqrt{N_{\alpha}}\right)/\left(x_{k}^{2}-N_{\alpha}\right)$, so
$\left|\omega_{k}-1\right|=\sqrt{\frac{4\left|N_{\alpha}\right|}{x_{k}^{2}-N_{\alpha}}}=2\sqrt{\frac{\left|N_{\alpha}\right|}{d}}\frac{1}{y_{k}}.$
Since $\left|\sqrt{\omega_{k}}+1\right|<2$, it follows that
$\left|\sqrt{\omega_{k}}-1\right|=\frac{\left|\omega_{k}-1\right|}{\left|\sqrt{\omega_{k}}+1\right|}>\sqrt{\frac{\left|N_{\alpha}\right|}{d}}\frac{1}{y_{k}}.$
@@
$\left|\frac{(\omega_{k}-1)^{2r_{0}+1}}{\left(1+\sqrt{\omega_{k}}\right)^{2r_{0}}}\right|=\left|\left(\omega_{k}-1\right)\left(\sqrt{\omega_{k}}-1\right)^{2r_{0}}\right|>2\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1}.$
Applying this inequality together with the upper bound for $\left|Y_{1,4,r_{0}}\left(\omega_{k}\right)\right|$ in (4.14) and the inequalities (2.3) in Lemma 2.3(b) to (4.3), it follows that
$\left|\omega_{k}^{1/4}-\zeta_{4}\frac{p}{q}\right|$ $>\frac{0.2915}{4^{r_{0}}\cdot r_{0}^{1/2}}\left(\frac{\left|N_{\alpha}\right|}{d}\right)^{r_{0}+1/2}\left(\frac{1}{y_{k}}\right)^{2r_{0}+1}.$
@@
since $N_{\alpha}$ is a square and $d<\left|N_{\alpha}\right|$. But this inequality is impossible. Hence this case cannot hold.
$4.$ $r_{0}>1$, $\zeta_{4}p/q\neq p_{r_{0}}/q_{r_{0}}$ for all $4$-th roots of unity, $\zeta_{4}$
@@
Since $N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b), so
@@ 4.5. $r_{0}>1$ and $\zeta_{4}p/q=p_{r_{0}}/q_{r_{0}}$ for some $4$-th root of unity, $\zeta_{4}$
Since $N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b). Also, as $r_{0}-1\geq 1$, it follows that $0.7405\cdot 0.0002344^{r_{0}-1}>0.0001735^{r_{0}-1}$. So the above inequality implies
@@ 4.6. Small $d$
We check directly all pairs $(a,d)$ of positive integers with $2\leq d\leq 104$ not a square and $N_{\alpha}=d-a^{2}$ a square. Note that there are only finitely many such pairs, since $N_{\alpha}=a^{2}-d<0$.
@@
Next we treat the case when K<1K<-1, where, as in Lemma 3.5, $K$ is the largest negative integer such that $y_{K}>1$.
@@
If K<1K<-1, then there are at most two distinct integer squares among the $y_{k}$’s.
@@
From Lemma 3.6, we know that K<1K<-1 can only happen in the following two cases:
@@
For the remainder of this section, we may assume that K=1K=-1.
@@
Similarly, if $N_{\alpha}$ is an odd square with $N_{\alpha}\leq-9$ and $u\geq 3$, then we have $y_{k}>(81/10)\left|N_{\alpha}\right|^{2}\geq(729/10)\left|N_{\alpha}\right|>(729/10)\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$. But, we saw at the start of this section that for $N_{\alpha}$ odd, there can be no squares $y_{k}$ and $y_{\ell}$ satisfying $y_{\ell}>y_{k}>\max\left(1,19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}\right)$. So we need only consider $u\leq 2$ when $N_{\alpha}$ is odd.
@@
For $u=1$, we proceed similarly. Here we have $y_{k}>\left|N_{\alpha}\right|^{2}/10$ for $|k|\geq 2$, so we need to consider $4\leq\left|N_{\alpha}\right|\leq 36$ when $N_{\alpha}$ is an even square and $9\leq\left|N_{\alpha}\right|\leq 169$ when $N_{\alpha}$ is an odd square.
@@
$y_{1}>\frac{57.32d^{2}}{\left|N_{\alpha}\right|^{2}}y_{-1}^{3}\geq\frac{57.32d^{2}}{\left|N_{\alpha}\right|^{2}}\left(\frac{\left|N_{\alpha}\right|u^{2}}{4}\right)^{3}=\frac{57.32d^{2}\left|N_{\alpha}\right|u^{6}}{64}.$
@@
Suppose that $N_{\alpha}$ is an odd square and $y_{\pm 1}$ is a square integer.
@@
(a) Next we show that if $t^{2}-du^{2}=4$, then $4\nmid t$. If $4\mid t$, then $du^{2}\equiv 12\pmod{16}$. Since $N_{\alpha}$ is an odd square, we have $N_{\alpha}\equiv 1,9\pmod{16}$. If $a$ is odd, then $a^{2}\equiv 1\pmod{8}$. So $d\equiv 2\pmod{8}$ and hence $u^{2}\equiv 6\pmod{8}$, which is not possible. If $a$ is even, then $a^{2}\equiv 0,4\pmod{16}$. Since $N_{\alpha}=d-a^{2}\equiv 1,9\pmod{16}$, it follows that $d\equiv 1\pmod{8}$. So $u^{2}\equiv 12\pmod{16}$, which is also impossible. So $4\nmid t$.
@@
Since $N_{\alpha}$ is odd, we have $N_{\alpha}\equiv 1,9\pmod{16}$. Also $a^{2}\equiv 0,1,4,9\pmod{16}$. Hence $d\equiv 1,2,5,9,10,13\pmod{16}$ and so $8|u^{2}$. Thus $4|u$.
@@
If $N_{\alpha}$ is an odd square, $1\leq u\leq 8$ and $y_{\pm 1}\in\mathbb{Z}$ a square with $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $u=2$, $t^{2}-du^{2}=-4$ and $\gcd\left(a^{2},d\right)=1$.
@@
From Lemma 5.4(a), if $t^{2}-du^{2}=4$ and $N_{\alpha}$ is an odd square, then $4|u$, so we are left with $u=4$ or $8$ here.
@@
Arguing modulo $9$, from $t^{2}-64d=4$, we see that $d\equiv 0,3,5,6\pmod{9}$. From this, $N_{\alpha}=d-a^{2}$ being a square and the squares modulo $9$ being $0,1,4,7$, we must have $\left(a^{2}\bmod 9,d\bmod 9\right)=(0,0)$, $(1,5)$, $(4,5)$ or $(7,5)$.
@@
Therefore by Lemma 3.12, we have $\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\geq 36$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}<(76/36)\left|N_{\alpha}\right|$. But by Lemma 3.5(c), we know that $y_{\pm 1}\geq 16\left|N_{\alpha}\right|$ here. So the case where $u=8$, $N_{\alpha}$ is an odd square and $t^{2}-du^{2}=4$ is excluded.
@@
We suppose otherwise. Since $N_{\alpha}$ is odd, $\gcd\left(a^{2},d\right)$ must be odd and since $d=\left(t^{2}+4\right)/4$, any odd prime factor, $p$, of the gcd must also be a factor of $t^{2}+4$. This means that $t^{2}\equiv-4\pmod{p}$. Hence $p\equiv 1\pmod{4}$. Since $p^{2}|a^{2}$ and $N_{\alpha}=d-a^{2}$ is a perfect square, this also implies that $p^{2}|d$. Therefore by Lemma 3.12, we have $\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\geq 100$. Hence the theorem holds if $y_{\pm 1}>(76/100)\left|N_{\alpha}\right|$, by Proposition 4.1. But by Lemma 3.5(c), we know that $y_{\pm 1}\geq\left|N_{\alpha}\right|$ here. So the theorem holds if $u=2$, $N_{\alpha}$ is an odd square and $\gcd\left(a^{2},d\right)>1$. ∎